OpenStudy (anonymous):

7a=2-b, 7a=5-b how do I us the substition method for this

jimthompson5910 (jim_thompson5910):

Hint: if 7a is both 2-b and 5-b, then 2-b = 5-b

OpenStudy (anonymous):

is that the method

jimthompson5910 (jim_thompson5910):

yes I'm using substitution

OpenStudy (anonymous):

oh thank you so much

jimthompson5910 (jim_thompson5910):

7a=5-b 2-b = 5-b ... substitute '2-b' in for '7a' (since they are the same)

OpenStudy (anonymous):

thanks I have 30 questins to answer and I am horible in algelbra.

jimthompson5910 (jim_thompson5910):

I'm sure you'll get better with more practice

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np